OpenStudy (thomas5267):

How do I combine this expression? \[ (6-t)^\frac{2}{3}-\frac{2t}{3}(6-t)^\frac{-1}{3} \]

OpenStudy (thomas5267):

Question edited. Please refresh.

OpenStudy (anonymous):

\[(6-t)^{\frac{2}{3}}-\frac{2t}{3(x-t)^{\frac{1}{3}}}\] now add like with fractions

OpenStudy (anonymous):

I think =1/(6-t)^(1/3)[6-t-2t/3] =1/(6-t)^(1/3)[(18-5t)/3]

OpenStudy (anonymous):

\[\frac{(6-t)^{\frac{2}{3}}\times3(6-t)^{\frac{1}{3}}-2t}{3(6-t)^{\frac{1}{3}}}\]

OpenStudy (thomas5267):

Then I am stuck at your step. My brain isn't functioning properly today... @satellite73

OpenStudy (anonymous):

\[\frac{(6-t)-2t}{3(6-t)^{\frac{1}{3}}}=\frac{6-3t}{3(6-t)^{\frac{1}{3}}}=\frac{2-t}{(6-t)^{\frac{1}{3}}}\]

OpenStudy (anonymous):

which step, one or two?

OpenStudy (thomas5267):

Step two, but isn't the third post incorrect. I think it should look like this. \[ \begin{align*} &\frac{(6-t)^{\frac{2}{3}}\times3(6-t)^{\frac{1}{3}}-2t}{3(6-t)^{\frac{1}{3}}} \\ =&\frac{3(6-t)-2t}{3(6-t)^\frac{1}{3}} \\ =&\frac{18-5t}{3(6-t)^\frac{1}{3}} \end{align*} \]

OpenStudy (anonymous):

yeah you got me. i forgot about that 3 in the numerator didn't i?

OpenStudy (anonymous):

step 2 is because \[a-\frac{b}{c}=\frac{ac-b}{c}\]